create new tag
view all tags

Shift Log 20111206

Laser focus referenced offline: -35650 to -35670 on manipulator vertical.

Background ~4700 with 20dB attenuation.


At meeting vertical emittance in extraction line reported to be between 18 - 35pm (OTR emittance strongly charge dependent).


Ext emittance shift extended to complete emittance measurements with wire scanners to try and confirm OTR measurements (need to check all wires are out of the beam line before starting).


Laserwire shift start. Signal attenuation reduced to 20dB. Timing search. Laser signal on APD ~ 45mV. Signals overlapped at 3.5V on the phase shifter.

Witnessed jumping between rf buckets. If under 3V laser and e separated. Started chamber at 60v, -310h, found rough middle 400v, 50h. BPM MFB2FF unsaturated in y, slightly saturated in x. Laser set at 100us, doing vertical chamber scan, 50 steps between 400 and 500. No signal - scanning between 300 and 600. No signal. Scanning between 600 and 800.


Laser timing unstable. Changed phase and overlapped pulses again.

Phase was 3.5V but Changing to 5V made much more stable again. Stability could be seen in lock box for laser which would occasionally jump from ~8ps noise to >>35ps. After the change to 5V this wasn't observed and the phase noise was around 3 - 4ps.

May check again later on.


Ambiguity in alignment of screen with respect to laser along electron beam path. This would in the end only affect vertical alignment.

Trying +- 200 microns in horizontal and then vertical scans at each location.

CV 400 - 600 CH: 80 P:5V - nothing
CV 400 - 600 CH: 280 P: 5V - nothing
CV 400 - 600 CH: 4 80 P: 5V - nothing
CV 400 - 600 CH: -100 P: 5V - nothing
CV 400 - 600 CH: -200 P: 5V - nothing
CV 300 - 700 CH: 80 P: 5V - nothing


Another timing check - reducing the OTR signal on the APD showed that the two signals were a little apart. Retuned to overlap.

Changed TD2 timing by 1 RF bucket of 357MHz. Phase shifter voltage now 2.5V.

CV 400 - 600 CH: 80 P: 2.5V - nothing

Going to redo screen alignment to beam and take good data during this - for information and for software development.

ManVer ManHor

Angle at IP based on CBPM readings at QM13 and 14FF are Y 100murad - thanks YoungIm!


YoungIM IPBPM background scan. IPBPM is 12mm wide and 6mm vertically

Position Log Time Stamp
V:1000, H:80 20111206_2204
V:-1000 H:80 20111206_2208
V:520 H:1000 20111206_2210
V:520 H:-1000 20111206_2212

No apparent change in background conditions. IPBPM doesn't make any difference. V:520 is beam centre position.


CV: 0 - 500 CH: 80 P: 2.50V - nothing
CV: 0 - -200 CH: 80 p 2.50V - nothing

Then realised the laser software interlock had not been turned off so these scans had no laser energy. Reset laser to 100us.

CV: -100 - 500 CH: 80 P: 2.50V - nothing

Checked the reference of the laser position to the screen - still between -35650 and -35670 on the vertical position.


Switched to OTR post IP line. CV: 400 CH:80 MV: -35660. Moved to CV 700. Image looks bright and of a hard edge.

Image file 20111206 folder OTRimage1.bit

Observed electron beam on OTR 23_39.fit At MV -35660 screen notch clearly visible on beam. Moved to MV -35760 and changed mirror angles to bring beam image into centre of field of view. Then changing lens position to bring image into focus (23_43_f_9mm.fit, 23_45_f_10mm.fit). Moved to MV -35860. Best focus seemed to be with f = 8mm. Moved back to MV -35660 to observe screen notch on image then moved CV to 680 (notch visible), 660 (notch visible, beam merged into one rather than two stripes), 600 (0002.fit)

Quad settings:

QM15FF 27.072A
QM14FF: -57.575A
QM13FF: 37.142A
QM12FF: 15.528A

At CV: 600 beam half obscured by notch in screen so we are not centred horizontally. Move to CH 280 (0004.fit), wrong way (beam more exposed). Move to CH -200 (0005.fit), slightly too far. Move to CH -150 (0007.fit). Move to CH 80 (0010). Move to CH 0 (0011).


Switch back to laserwire post IP setting to try laserwire scanning again. Regen timing changed so timing redone.

CV: 500 - 700 CH: 0 P: 7.5V - nothing
CV: -100 - 800 CH: 0 P: 7.5V - nothing


Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r2 - 06 Dec 2011 - LaurieNevay

This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2024 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback